
None.
------ 2nd half:
- Pointing offsets larger than previous nights (~7 arcmin).
- Mirror started getting warmer from about 3:00 UTC - did the chiller stop working?
- Observer2 computer really sluggish.
- Observer Console was behaving oddly at least for exposure 174804: the blue exposure bar didn't start, instead the number in the grey bar kept fluctuating and gradually decreasing. The exposure was actually being taken, as also indicated by a second instance of Observer Console opened on the Observer2 computer.
- Lost guide stars once; guide star display afterwards doesn't update to "live view"
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2013-02-02 | 8.5 | 0.0 | 0.0 | 0.0 | 0.0 | 8.5 |
Links:
[1] mailto:frieman@fnal.gov
[2] mailto:anja@slac.stanford.edu
None.
------- 2nd half:
- Sky backgrounds in z-band higher than yesterday, despite less moonlight. Rasicam shows more water vapor than in previous nights, so presumably this is OH emission?